OpenStudy (anonymous):

VERY confused: f(x)=3x g(x)=x-3 and h(x)= vTx What is (fogoh)(25) isn't it 3(x-3) 3x-3 ? Then I don't know what to do!

OpenStudy (anonymous):

h = what?

OpenStudy (anonymous):

Oh, that's how I write the square root of x haha, sorry.

OpenStudy (anonymous):

so take it step by step... what is the innermost fucntion? what's next innermost? what's the outermost? \[\left( f \circ g \circ h \right)\left( x \right)=f \left( g \left( h \left( x \right) \right) \right)\] start plugging in the innermost...\[\left( f \circ g \circ h \right)\left( 25 \right)=f \left( g \left( h \left( 25 \right) \right) \right)\]

OpenStudy (anonymous):

okay, then I plug in h right?

OpenStudy (anonymous):

\[f(g (\sqrt{x}(25)))\]

OpenStudy (anonymous):

close... \[f \left( g \left( \sqrt{25} \right) \right)\]

OpenStudy (anonymous):

remember, x = 25

OpenStudy (anonymous):

Okay now G, correct?

OpenStudy (anonymous):

\[\sqrt{25}-3\]

OpenStudy (anonymous):

so what is h(25)? it's best to simplify as you go... you should also see if you can come up with (f o g o h)(x) by itself and then plug in x = 25 to check that you get the same number.

OpenStudy (anonymous):

Oh it's 5 haha, whoops.

OpenStudy (anonymous):

so do I do 5-3, and then multiply that by 3?

OpenStudy (anonymous):

you know it!

OpenStudy (anonymous):

good luck, got to go!